OpenStudy (anonymous):

Choose one of the factors of x^3 – 1331 x – 11 x^2 – 11x + 121 x^2 + 22x + 121 None of the above

OpenStudy (anonymous):

Hi

OpenStudy (radar):

The difference of two perfect cubes x and 11

OpenStudy (anonymous):

yes i think that the answer would be x-11

OpenStudy (radar):

Let me check:(x-11)(x^2+11x+121)

OpenStudy (radar):

It is x-11 .

OpenStudy (anonymous):

yay do you think you can help me with one more?

OpenStudy (radar):

Just one lol

OpenStudy (anonymous):

yes one more i promise.

OpenStudy (radar):

go

OpenStudy (anonymous):

Choose one of the factors of 5x3 – 135

OpenStudy (anonymous):

its 5x^3-135

OpenStudy (radar):

5(x3-27)

OpenStudy (anonymous):

ok i got that part

OpenStudy (radar):

Now notice that we x3-27 which is the difference of two perfect cubes, that can be factored further.

OpenStudy (anonymous):

and the cubic root of 27 would be 3 right?

OpenStudy (radar):

(x-3)(x2 +3x+9)(5) what are your choices? Yes

OpenStudy (anonymous):

5 x – 3 x2 + 3x + 9 All of the above

OpenStudy (radar):

What do you think is the best answer?

OpenStudy (anonymous):

all of them

OpenStudy (radar):

Bingo!

OpenStudy (anonymous):

also can you explain how you got, (x-3)(x2 +3x+9)(5) please.

OpenStudy (radar):

The difference of two cubes, is a special product that you need to memorize or know where to look. \[(a ^{3}-b ^{3})= (a-b)(a ^{2}+ab + b ^{2)}\]

OpenStudy (radar):

The sum of two cubes is also a special case and has a similar product

OpenStudy (anonymous):

Is that a special formula to know where to plug the numbers in?

OpenStudy (radar):

Yes, a was x and b was 3

OpenStudy (anonymous):

so ok to be good with this, (a3−b3)=(a−b)(a2+ab+b2) would be substituted by (x-3)(x^2+3x+3^2)

OpenStudy (anonymous):

(5) at the end so if we check our work and multiply it by 5 we would get the same thing in the beginning.

OpenStudy (radar):

Yes and if it was (a3+b3) it would factor to: (a+b)(a2-ab+b2)

OpenStudy (anonymous):

alright thank you very much. I have a test tomorrow so now I'm ready.

OpenStudy (radar):

Good luck on the test.

OpenStudy (anonymous):

thanks